OpenStudy (anonymous):

Medal Solve each system of equations by addition-subtraction, or by substitution. 4n = 27 -3m m=11-2n

OpenStudy (danjs):

they have one already solved for m for you, just put that value in for m in the other one

OpenStudy (danjs):

4n=27-3*(11-2n) n = ... If n is that, you can figure m from either equation

OpenStudy (anonymous):

4n = 27 - 3 ( 9n)

OpenStudy (anonymous):

@DanJS am i right?

OpenStudy (danjs):

no, not sure what you did distribute the 3 into the quantity, then do additions

OpenStudy (danjs):

4n = 27 + (-3)*(11-2n) 4n = 27 - 33 + 6n

OpenStudy (anonymous):

n= -30

OpenStudy (anonymous):

m i right i solved it through

OpenStudy (danjs):

4n-6n = -6 -2n = -6 n = 3

OpenStudy (anonymous):

i added 27 by 33 thats my mistake

OpenStudy (danjs):

you would have to show your steps, not sure what you did

OpenStudy (danjs):

ahh

OpenStudy (anonymous):

to find m i sub in 3

OpenStudy (danjs):

yeah distribuiting the quantity gives a - 33

OpenStudy (danjs):

yeah, you know one of the 2 variables, you can use any relationship between the two to figure the other one

OpenStudy (anonymous):

i got m = 5

OpenStudy (anonymous):

4 (3) = 27 = 3m 12 = 27 - 3m -15 = -3m m = 5

OpenStudy (danjs):

yes, to check if you are correct, try the value in the other equation , should get the same answer

OpenStudy (anonymous):

thank you so much